ebhancoc:

A rabbit's fur grows at the rate of 5/2 inches per month. How many months will it take the rabbit's fur to grow 15/4 inches? Explain your answer using words, and show your work using division of fractions.

jhonyy9:

this is the right way so bc. you need getting for rabbit 15/4 inches to grow and you know that the rabbit grows at the rate of 5/2 inches per month so divide this 15/4 by 5/2 and yill get the right answer

jhonyy9:

can you divide it ?

jhonyy9:

(15/4)/(5/2) = (15/4)×(2/5) = ?

jhonyy9:

can you multiplie these fractions ?

jhonyy9:

= (15×2)/(4×5) = ?

jhonyy9:

courage i m here to help you

jhonyy9:

15×2 = ?

jhonyy9:

please collaborate

ebhancoc:

15/4 by 5/2 = 1.5

jhonyy9:

yes sure so what mean this ?

jhonyy9:

so this mean that the rabbit need one month and a half month to grow to 15/4 inches ok. ?

jhonyy9:

do you understand it now ?